OpenStudy (anonymous):

There is no solution to the system below. 2x + 4y = 8 x + 2y = 6 true or false?

OpenStudy (anonymous):

x=6-2y 2(6-2y)+4y=8 12-4y+4y=8 12=8 no solutions

OpenStudy (anonymous):

2x+4y=8 x+2y=6/*2 2x+4y=8 2x+4y=12 0=-4 no solutions

OpenStudy (anonymous):

yeah right no solution